If the probability data will fall between the mean and a score above the mean is 0.341 and the probability data will fall between the mean and a score below the mean is 0.477, what is the probability the data will fall between those two scores above and below the mean?
Answer #1

P(between mean and mean + a) = 0.341

P(between mean - b and mean) = 0.477

P(between mean - b and mean + a) = P(between mean and mean + a) + P(between mean - b and mean) = 0.341 + 0.477 = 0.818